find the total no. of numbers less then 1000 & divisible by 5 that can be formed with the digits 0,1,2,3,4,5,6,7,8,9(none of the digits repeated).
Answer by sudhanshu_kmr(1152)


less then 1000 means upto 3 digit numbers and at unit place either 0 or 5.

Single digit numbers...... only 5 i.e no. of ways = 1

Two digit numbers...
when 0 on unit place , no. of ways = 9

when 5 on unit place, no. of ways = 8

total no. of ways for 2 digit = 17

Three digit numbers..

when 0 on unit place , no. of ways 9P2 = 9*8 = 72

when 5 on unit place, no. of ways = 9P2 - 8 = 64
( in 8 numbers 0 will be on first place, so discard it)
total no. of ways for 3 digit = 136



total no. of ways = 1 + 17 + 136 = 154